# NovelQA: A Benchmark for Long-Range Novel Question Answering

### v1

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has introduced a new frontier in natural language processing, particularly in understanding and processing long-context information. However, the evaluation of these models’ long-context abilities remains a challenge due to the limitations of current benchmarks. To address this gap, we introduce NovelQA, a benchmark specifically designed to test the capabilities of LLMs with extended texts. Constructed from English novels, NovelQA offers a unique blend of complexity, length, and narrative coherence, making it an ideal tool for assessing deep textual understanding in LLMs. This paper presents the design and construction of NovelQA, highlighting its manual annotation, and diverse question types. Our evaluation of Long-context LLMs on NovelQA reveals significant insights into the models’ performance, particularly emphasizing the challenges they face with multi-hop reasoning, detail-oriented questions, and extremely long input with more than 100,000 tokens. The results underscore the necessity for further advancements in LLMs to improve their long-context comprehension and computational literary studies.
NovelQA: A Benchmark for Long-Range Novel Question Answering
Cunxiang Wang1111,2*2*2 *, Ruoxi Ning2222thanks:   Equal Contribution, Boqi Pan3333, Tonghui Wu3333, Qipeng Guo4444,
Cheng Deng5555, Guangsheng Bao1111,$2$ Qian Wang3333 and Yue Zhang1111thanks:    The correponding author. $1$School of Engineering, Westlake University; $2$Zhejiang University $3$Hangzhou Normal University; $4$Shanghai AI Lab; $5$SJTU {wangcunxiang, zhangyue}@westlake.edu.cn; ruoxining@outlook.com
@@ 1 Introduction
Recent years have seen a remarkable surge in the development of Large Language Models (LLMs) (OpenAI, 2023a; Touvron et al., 2023). Among these developments, long-context LLMs stand out for their ability to process and interpret extended pieces of text (Tworkowski et al., 2023; Team, 2023; Anthropic, 2023). This capability is essential for complex tasks that require a deep and nuanced understanding of lengthy documents, such as legal cases Xiao et al. (2021) or academic papers Groeneveld et al. (2024), where the key is to understand extended narratives Xu et al. (2023); Wang et al. (2023d). Meanwhile, the ability to analyze multiple long documents simultaneously is increasingly crucial, supporting more informed decision-making in various fields Deng et al. (2023); Lin et al. (2023); Boiko et al. (2023).
@@
To fill this gap, we introduce NovelQA, a benchmark crafted to specifically evaluate LLMs’ performance on texts with context windows exceeding 100,000 tokens. Unlike existing benchmarks (Shaham et al., 2023; An et al., 2023; Adams et al., 2024), NovelQA addresses the need for assessing extremely long-context understanding, offering a refined and comprehensive tool for advancing natural language processing capabilities.
Figure 2: Illustrative Examples from NovelQA: This figure showcases two sample questions. For each question, models are evaluated under two distinct settings – Multichoice, where the task is to select the correct answer from four options, and Generative, where the model generates an answer.
@@
We construct NovelQA based on novels in English. In addition to their value in computaitonal literature ad creativity studies, novels are also ideal for testing long-context modeling because they are long and complex, with plots that are closely linked from start to end. We select novels from various eras, genres, and formats to enhance diversity. The annotation process is performed by a group of skilled annotators, all of whom are holding or pursuing a degree in English Literature and have a strong interest in and familiarity with the novels they annotate. Each question is paired with a ‘golden answer’ and corresponding textual evidences from the novels. And we categorize them by complexity and aspect for detailed analysis. Figure 2 presents two examples, while Table 8 details the distribution across types, with multi-hop, single-hop, and detailed questions constituting 35.0%, 42.8%, and 22.2% of the dataset, respectively.
@@
In our NovelQA experiments, we assess various long-context LLMs, including commercial models such as GPT-4-128K (OpenAI, 2023b) and Claude 2.1-200K (Anthropic, 2023), alongside open-source models like InternLM2-Chat (Team, 2023). The primary findings offer a significant insight: even the most advanced long-context LLMs face challenges in consistently extracting and processing accurate information from extended texts. For example, GPT-4, the top performer, achieves a 46.88% accuracy rate, whereas the open-source InternLM2-20b achieves 32.37% in a generative setting. This difficulty is particularly apparent in answering multi-hop questions and queries that probe meanings, relationships, spans, and timelines, highlighting a significant gap in the models’ long-range comprehension. Moreover, our data shows a decline in performance for evidence situated beyond the 100,000-token mark, including information at the novel’s end, diverging from the anticipated lost-in-middle phenomenon (Liu et al., 2023). This shift suggests a distinct challenge faced by LLMs when processing texts exceeding 100k tokens in length. These results highlight challenges not only in memory optimization but also in the nuanced comprehension and integration of lengthy texts, indicating a substantial obstacle on the path to truly effective long-context LLMs.
@@
To the best of our knowledge, NovelQA is the first long-context QA benchmark featuring manually crafted questions, golden answers, and evidences, with contexts extending beyond 200,000 tokens. It also is the first long-range QA dataset within computational literary studies.  [^1] We have released the demonstrations and input of the testset, and created a leaderboard. More details can be found in https://github.com/NovelQA/novelqa.github.io.
@@ 2 Related Work
Long-Range Benchmarks Evaluating the ability of Long-Context Pretrained Language Models has been a hot topic (Tay et al., 2021; Shaham et al., 2023; Pang et al., 2022; Wang et al., 2024b, 2023b; Yu et al., 2024). When entering the era of Large Language Models, the context window length has been much longer than ever, and the decoder-only LLMs have been the mainstream of language models and they have faced specific problems, such as the Lost-in-middle issue (Liu et al., 2023). Thus, increasing benchmarks are created for evaluating long-context LLMs. Among those benchmarks, several are representatives. ZeroSCROLLS (Shaham et al., 2023) and LooGLE (Li et al., 2023a) both emphasize the importance of understanding and aggregating information from long texts, presenting challenges that highlight areas for improvement in current models. L-Eval (An et al., 2023) introduces a suite of tasks with human-labeled query-response pairs to assess LLMs’ performance in processing long inputs effectively, using advanced metrics for a more accurate evaluation. Furthermore, benchmarks like LongBench (Bai et al., 2023), BAMBOO (Dong et al., 2023a), and LongBench-Chat (Bai et al., 2024) offer a diverse set of tasks across languages and domains, from reasoning and coding to summarization and multilingual translation. These benchmarks are designed to rigorously test the ability of LLMs to manage extensive contexts, with LongBench-Chat specifically focusing on instruction-following capabilities in long-context interactions. Additionally, LongHealth (Adams et al., 2024) addresses the need for LLMs to interpret long clinical documents accurately, providing a specialized benchmark for evaluating models on medical texts. This focus on domain-specific challenges underscores the broader necessity for LLMs to not only handle long texts but to do so in a manner that is accurate and contextually relevant across various fields.
@@
NovelQA distinguishes from existing benchmarks in three points. Firstly, NovelQA features an average token length exceeding 200,000, far surpassing the tens of thousands typically found in other benchmarks. Secondly, while other datasets often rely on LLMs or existing datasets for content creation, the questions, golden answers, and evidences of NovelQA are entirely crafted through human effort. Lastly, it is the first QA dataset specifically designed for the literature domain.
Long-Context Language Modeling can be divided into several parts including efficient attention, preserving long-term memory like using KV cache, extrapolative positional embedding module, context pre/post-processing Huang et al. (2023); Dong et al. (2023b). Using efficient attention can reduce computational complexity and memory usage Beltagy et al. (2020); Wang et al. (2020); Tworkowski et al. (2023); Li et al. (2023b). Preserving KV cache or context-level cache allows models to recall and leverage past information without reprocessing, enhancing coherence over long texts Chevalier et al. (2023); Wu et al. (2022a, b); Lin et al. (2024); Zhong et al. (2023); Chen et al. (2023a); Hooper et al. (2024). Deploying an extrapolative positional embedding to extend beyond the sequence lengths seen during training Chen et al. (2023b); Su et al. (2024). Pre/post-processing the context can make models focus on key information Li et al. (2023c); Jiang et al. (2023); Zhang et al. (2024); Han et al. (2023). Moreover, some training methods are put forward to support further long-context language modeling development Wang et al. (2024a); Press et al. (2021). Since these models are more about exploring long-context language modeling without scaling in data and model size, the focus is on language modeling indicators, such as BLEU and Perplexity, NovelQA can serve as a benchmark for evaluating those long-context language modeling methods.
Computational Literary Studies in Novels as an inseparable part of the literature field, are a notable theme in NLP. Past research on novels has covered topics such as word segmentation and POS-tagging (Zhang et al., 2014; Liu and Zhang, 2012; Qiu and Zhang, 2015) and event extraction (Makazhanov et al., 2014; Sims and Bamman, 2020). Previous models’ limited input window size hindered a comprehensive examination of long novels. With the advancement of models with long-context capabilities, research has been launched on benchmarks of novel summarization (Kryściński et al., 2021; Chang et al., 2023b) and evaluation on LLM’s inherent novel knowledge (Chang et al., 2023a). NovelQA pioneers in question-answering on lengthy novels, marking a contribution to novel studies in the NLP field.
@@ 3.1 Dataset Description
Data Formulation Every novel ($N$) in the dataset corresponds to multiple pieces of annotated data ($d_{i}$). Each piece of data consists of the following domains, question ($Q_{i}$), answer ($A_{i}$), multi-choices ($a_{i,0}$, $a_{i,1}$, $a_{i,2}$, and $a_{i,3}$), gold label ($a_{i,gold}$), evidences ($s_{i,0}$, $s_{i,1}$, …), and type ($Complx_{i}$ and $Aspect_{i}$), among which the answer domain is a short answer to the question, while the multi-choices and gold label domain forms multi-choices, indicating that our dataset can serve to the evaluation of both the generative task and the multichoice task. In the generative setting, a novel text $N$ and a question $Q_{i}$ are combined to send into the model each time, and the generated answer is compared with the answer $A_{i}$. In the multichoice setting, the novel $N$, a question $Q_{i}$, and the four choices $a_{i,0}$ to $a_{i,3}$ are sent into the model, and the output is evaluated according to the gold label $a_{i,gold}$. Meanwhile, the evidences domain consists of either the original excerpts from the novel, or the reasoning steps written by the annotator.
@@
Book Selection We aim to enhance diversity by selecting novels from various eras, genres, and formats. Recognizing that newer, popular novels are still under copyright protection, NovelQA inevitably includes some restricted-access works. Our efforts to respect copyright laws are detailed in the Ethical Statements section. We source open-copyright novels from Project Gutenberg [^2] https://www.gutenberg.org/, and purchase e-books.All selected books exceed 50k words (approximately 67k tokens) and are in English.A visualization of the distribution of book token lengths is provided in Appendix A.2.

### v2

# NovelQA: Benchmarking Question Answering on Documents Exceeding 200K Tokens

@@ Abstract
The rapid advancement of Large Language Models (LLMs) has introduced a new frontier in natural language processing, particularly in understanding and processing long-context information. However, the evaluation of these models’ long-context abilities remains a challenge due to the limitations of current benchmarks. To address this gap, we introduce NovelQA, a benchmark specifically designed to test the capabilities of LLMs with extended texts. Constructed from English novels, NovelQA offers a unique blend of complexity, length, and narrative coherence, making it an ideal tool for assessing deep textual understanding in LLMs. This paper presents the design and construction of NovelQA, highlighting its manual annotation, and diverse question types. Our evaluation of Long-context LLMs on NovelQA reveals significant insights into the models’ performance, particularly emphasizing the challenges they face with multi-hop reasoning, detail-oriented questions, and extremely long input with an average length more than 200,000 tokens. The results underscore the necessity for further advancements in LLMs to improve their long-context comprehension and computational literary studies.
@@ 1 Introduction
Recent years have seen a remarkable surge in the development of Large Language Models (LLMs) (OpenAI, 2023a; Touvron et al., 2023). Among these developments, long-context LLMs stand out for their ability to process and interpret extended pieces of text (Tworkowski et al., 2023; Team, 2023; Anthropic, 2023). This capability is essential for complex tasks that require a deep and nuanced understanding of lengthy documents, such as legal cases Xiao et al. (2021) or academic papers Groeneveld et al. (2024), where the key is to understand extended narratives Xu et al. (2023). Meanwhile, the ability to analyze extremely long documents and multiple documents simultaneously is increasingly crucial, supporting more informed decision-making in various fields Deng et al. (2023); Lin et al. (2023); Boiko et al. (2023).
@@
To fill this gap, we introduce NovelQA, a benchmark crafted to specifically evaluate LLMs’ performance on texts with context windows exceeding 200,000 tokens. Unlike existing benchmarks (Shaham et al., 2023; An et al., 2023; Adams et al., 2024), NovelQA addresses the need for assessing extremely long-context understanding, offering a refined and comprehensive tool for advancing natural language processing capabilities. We construct NovelQA based on novels in English, which are also ideal for testing long-context modeling because they are long and complex, with plots that are closely linked from start to end. We select novels from various eras, genres, and formats to enhance diversity. The annotation process is performed by a group of expert annotators, all of whom are holding or pursuing a degree in English Literature and have a strong interest in and familiarity with the novels they annotate. Each question is paired with a ‘golden answer’ and corresponding textual evidences from the novels. And we categorize them by complexity and aspect for detailed analysis. Figure 2 presents two examples, while Table 6 details the distribution of question types. The dataset includes multi-hop, single-hop, and detail questions, which test the model’s abilities to retrieve and integrate scattered information, retrieve information and summarize, and precisely identify specific and subtle details, respectively. The above capabilities are more challenging in a particularly long context. For example, finding relevant information is a particularly big challenge.
Figure 2: Illustrative Examples from NovelQA: This figure showcases two sample questions. For each question, models are evaluated under two distinct settings – multi-choice, where the task is to select the correct answer from four options, and Generative, where the model generates an answer.
@@
Using the dataset, we assess various long-context LLMs, including commercial models such as GPT-4-128K (OpenAI, 2023b) and Claude 2.1-200K (Anthropic, 2023), alongside open-source models like InternLM2-Chat (Team, 2023). The primary findings offer a significant insight: even the most advanced long-context LLMs face challenges in consistently extracting and processing accurate information from extended texts. For example, GPT-4, the top performer, achieves a 46.88% accuracy rate, whereas the open-source InternLM2-20b achieves 32.37% in a generative setting. In addition, some models are unable to answer properly when the context exceeds 100K tokens, even if their context windows are much longer (xverse, 2023; 01-ai, 2023). Moreover, operating LLMs on inputs exceeding 200,000 tokens presents technical challenges, particularly regarding memory requirements and associated costs. The difficulty of valid models is particularly apparent in answering multi-hop questions and queries that probe meanings, relationships, spans, and timelines, highlighting a significant gap in the models’ long-range comprehension. Moreover, our data shows a decline in performance for evidence situated beyond the 100,000-token mark, including information at the novel’s end, diverging from the anticipated lost-in-middle phenomenon (Liu et al., 2023). This shift suggests a distinct challenge faced by LLMs when processing texts exceeding 100K tokens in length. These results highlight challenges not only in memory optimization but also in the nuanced comprehension and integration of lengthy texts, indicating a substantial obstacle on the path to truly effective long-context LLMs.
@@
To the best of our knowledge, NovelQA is the first long-context QA benchmark featuring manually crafted questions, golden answers, and evidences, with contexts extending beyond 200,000 tokens.  [^1] We have released the demonstrations and input of the testset, and created a leaderboard. More details can be found in https://github.com/NovelQA/novelqa.github.io. And NovelQA is released under the Apache-2.0 License.
@@ 2 Related Work
Long-Range Benchmarks Evaluating the ability of Long-Context Pretrained Language Models has been a hot topic (Tay et al., 2021; Shaham et al., 2023; Pang et al., 2022; Wang et al., 2024a; Yu et al., 2024). When entering the era of Large Language Models, the context window length has been much longer than ever, and the decoder-only LLMs have been the mainstream of language models and they have faced specific problems, such as the Lost-in-middle issue (Liu et al., 2023). Thus, increasing benchmarks are created for evaluating long-context LLMs. Among those benchmarks, several are representatives. The Needle in A Haystack test (gkamradt et al., 2023) involves inserting an unrelated sentence into a long context and asking the model to retrieve it. However, this test does not represent a natural language question answering task that requires composite and complex capabilities for processing long contexts. ZeroSCROLLS (Shaham et al., 2023), LooGLE (Li et al., 2023a), Counting-Stars(Song et al., 2024), and LongICLBench(Li et al., 2024) emphasize the importance of understanding and aggregating information from long texts, presenting challenges that highlight areas for improvement in current models. L-Eval (An et al., 2023) introduces a suite of tasks with human-labeled query-response pairs to assess LLMs’ performance in processing long inputs effectively, using advanced metrics for a more accurate evaluation. Furthermore, benchmarks like LongBench (Bai et al., 2023), BAMBOO (Dong et al., 2023a), LongBench-Chat (Bai et al., 2024) offer a diverse set of tasks across languages and domains, from reasoning and coding to summarization and multilingual translation. These benchmarks are designed to rigorously test the ability of LLMs to manage extensive contexts, with LongBench-Chat specifically focusing on instruction-following capabilities in long-context interactions. Additionally, LongHealth (Adams et al., 2024) addresses the need for LLMs to interpret long clinical documents accurately, providing a specialized benchmark for evaluating models on medical texts. This focus on domain-specific challenges underscores the broader necessity for LLMs to not only handle long texts but to do so in a manner that is accurate and contextually relevant across various fields.
@@
NovelQA distinguishes from existing benchmarks in three points. Firstly, NovelQA features an average token length exceeding 200,000, far surpassing the tens of thousands typically found in other benchmarks. Secondly, while other datasets often rely on LLMs or existing datasets for content creation, the questions, golden answers, and evidences of NovelQA are entirely crafted through human effort. We present a detailed comparison between NovelQA and other benchmarks in Appendix Fig 4.
We also discuss related Long-Context Language Modeling methods in Appendix Sec 4.
@@ 3.1 Dataset Description
Data Formulation Every novel ($N$) in the dataset corresponds to multiple pieces of annotated data ($d_{i}$). Each piece of data consists of the following domains, question ($Q_{i}$), answer ($A_{i}$), multi-choices ($a_{i,0}$, $a_{i,1}$, $a_{i,2}$, and $a_{i,3}$), gold label ($a_{i,gold}$), evidences ($s_{i,0}$, $s_{i,1}$, …), and type ($Complx_{i}$ and $Aspect_{i}$), among which the answer domain is a short answer to the question, while the multi-choices and gold label domain forms multi-choices, indicating that our dataset can serve to the evaluation of both the generative task and the multi-choice task. In the generative setting, a novel text $N$ and a question $Q_{i}$ are combined to send into the model each time, and the generated answer is compared with the answer $A_{i}$. In the multi-choice setting, the novel $N$, a question $Q_{i}$, and the four choices $a_{i,0}$ to $a_{i,3}$ are sent into the model, and the output is evaluated according to the gold label $a_{i,gold}$. Meanwhile, the evidences domain consists of either the original excerpts from the novel, or the reasoning steps written by the annotator.
@@
Book Selection We source public domain novels from Project Gutenberg, and purchase e-books from internet if it is necessary. We aim to enhance diversity by selecting novels from various eras, genres, and formats. Recognizing that some newer, popular novels are still under copyright protection, NovelQA inevitably includes some copyrighted novels, comprising a mixture of 65 novels in public domain and 24 ones in copyright-protection.  [^2] We evaluate all annotated questions in this paper and conduct analysis. For the public access, we have released all constructed data. However, we only release public domain novels. Therefore, we offer two types of metrics in the evaluation system and leaderboard: one for evaluating QAs within public domain novels and another for evaluating QAs across all novels. All selected books exceed 50K words (approximately 67k tokens) and are in English. A visualization of the distribution of book token lengths is provided in Appendix Sec D.2.